OpenStudy (anonymous):

x^2-6x+9m give m for the equation two positive solutions

OpenStudy (anonymous):

are you just supposed to make up m?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok. I'll let m=1, then we have x^2-6x+9 factoring it: (x-3)^2=0 Factors are x=3 (with a multiplicity of two).

OpenStudy (anonymous):

ok ... this is one than let m=0

OpenStudy (anonymous):

ok, what do you get for m=0?

OpenStudy (anonymous):

x^2-6x

OpenStudy (anonymous):

ok, now factor out an x and set the whole thing equal to zero

OpenStudy (anonymous):

solution became 0 and 6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i think in quadratic x1*x2=c/a that's mean 9m >0 m>0 and Delta -36m+36 >0 mean m > 1 that's mean that the solution is from 0 to 1

OpenStudy (anonymous):

well, the quadratic formula will tell you the possible roots, yeah. first off, the discriminant must be larger than or equal to zero. We get,\[36-4(1)(9m)\ge0\]This gives, \[36\ge36m\] or,\[m \le1\] So if m is smaller than or equal to 1, we get real roots from this quadratic. If they are to be both positive then,\[\frac{6-\sqrt{36-36m}}{2}>0\]This gives,\[6-\sqrt{36-36m}>0\]and,\[36 > 36-36m\]and,\[m>0\]This m must be BOTH greater than 0 and less than or equal to 1. Thus m is in the interval (0,1], as you said :)

OpenStudy (anonymous):

Fine thank you ....

OpenStudy (anonymous):

No medal for me? lol